# Covariance of Scattering Amplitudes from Counting Carefully

### v2

@@ 3.4 Refined Counting Function
The exact counting function can be identified through the use of generating functionals. To that end note that polynomials can be used to count by assigning a power to each possible outcome up to some $n$, for example a six sided die can be represented by the polynomial
@@
Consider a set of variables $\vec{a}_{i}=(a_{i,0},\dots,a_{i,i})$ with $a_{i,0}$ corresponding to a vertex with $i$ legs and all being on-shell, while $a_{i,i}$ refers to the same vertex but with all legs off-shell. To remove the off-set, $\vec{a}_{1}$ will now be identified with the 1 point vertex, $\vec{a}_{2}$ the two point and so on.
@@ 4 On-Shell Connected Functions and Covariant Feynman Rules
where the sum over $S_{n}$ is a sum over all permutations of the external legs. It remains to show that replacing the coefficients $a_{i+2,l}$ with the individual Feynman rules $\delta^{i+2}\Gamma/\delta\phi^{i+2}$, applying the vacuum condition to all legs and the on-shell condition to the i+2li+2-l external legs leads to a series cancellations proving the covariance of $\mathcal{A}_{n}$.
@@ 4.2 Existence of Covariant Feynman Rules
The formula was derived by assuming that the different index contractions are only relevant for the non-tensor parts of the Feynman rules, this is sufficient for showing covariance and leads to the closed formula for the on-shell amputated function $\mathcal{A}_{a_{1}\dots a_{n}}$ provided in  alminawi2025scalaramps
@@
where $\mathcal{F}_{n,k}(\mathcal{R}_{3},\dots)$ are the polynomials counting Feynman diagrams given by (26) and $\sum_{S_{n}}$ represents a sum over all the permutations of the indices of $\mathcal{A}_{n}$. The placement of the indices is straightforward, with the only restriction being that the pair of indices in a propagator may not both be contracted in the same vertex; as that would correspond to a loop. To illustrate the use of the formula $\mathcal{A}_{6}$ is given by
@@ 5 Conclusion and Outlook
The methods used at the tree level lay out a simple path towards extending the results to arbitrary loop order by adjusting the counting function, which is planned for a future work, but it has been explored for 1 and 2 point functions at the 1-loop level in  Apostolos2020:Covariance ; alminawi2025scalaramps .

### v3

@@ 3.4 Refined Counting Function
The exact counting function can be identified through the use of generating functionals. To that end note that polynomials can be used to count by assigning a power and numerical coefficient to each possible outcome up to some $n$
$P_{n}(x)=\alpha_{0}+\alpha_{1}x+\alpha_{2}x^{2}+\alpha_{3}x^{3}+\dots+x^{n}=\sum_{i=0}^{n}\alpha_{i}x^{i}$ (25)
for example a six sided die can be represented by the polynomial
@@
Consider a set of variables $\vec{a}_{i}=(a_{i,0},\dots,a_{i,i})$ with $a_{i,0}$ corresponding to a vertex with $i$ legs and all being on-shell, while $a_{i,i}$ refers to the same vertex but with all legs off-shell. Since the focus of this counting is specific to Feynman diagrams, it makes sense to remove the off-set present in the graph theory conventions, thus $\vec{a}_{1}$ will now be identified with the 1 point vertex, $\vec{a}_{2}$ the two point and so on, thus the variables $a_{i}$ will correspond to vertices with $i$ legs, unlike in 3.3, where they corresponded to vertices with $i+2$ legs.
@@ 4 On-Shell Connected Functions and Covariant Feynman Rules
where the sum over $S_{n}$ is a sum over all permutations of the external legs. It remains to show that replacing the coefficients $a_{i+2,l}$ with the individual Feynman rules $\delta^{i+2}\Gamma/\delta\phi^{i+2}$, applying the vacuum condition to all legs and the on-shell condition to the i+2li+2-l external legs leads to a series cancellations proving that the transformation $\mathcal{A}_{n}\to\mathcal{A}^{\prime}_{n}$ under a field redefinition $\phi\to\psi(\phi)$ is given by
$\mathcal{A}_{n}=\left(\frac{\delta\psi}{\delta\phi}\right)^{n}\mathcal{A}^{\prime}_{n}$ (32)
@@
accordingly the covariance condition now reads
$\mathcal{A}_{n}=z_{1}^{n}\mathcal{A}^{\prime}_{n}$ (34)
@@ 4.2 Existence of Covariant Feynman Rules
The formula was derived by assuming that the different index contractions are only relevant for the non-tensor parts of the Feynman rules, this is sufficient for showing covariance and leads to the closed formula for the on-shell amputated function $\mathcal{A}_{a_{1}\dots a_{n}}$ provided in Alminawi et al. (2025)
@@
where $\mathcal{F}_{n,k}(\mathcal{R}_{3},\dots)$ are the polynomials counting Feynman diagrams given by (30) and $\sum_{S_{n}}$ represents a sum over all the permutations of the indices of $\mathcal{A}_{n}$. The placement of the indices is straightforward, with the only restriction being that the pair of indices in a propagator may not both be contracted in the same vertex, nor may two vertices be contracted to each other more than once; as that would correspond to a loop. To illustrate the use of the formula $\mathcal{A}_{6}$ is given by
@@
An explicit derivation of the covariant rules is easily achieved through the use of geometry as done in  Alminawi et al. (2025); Finn et al. (2020). However, a derivation that does not rely on geometric methods has not yet been identified and is slated for future works.
@@ 5 Conclusion and Outlook
The methods used at the tree level lay out a simple path towards extending the results to arbitrary loop order by adjusting the counting function, which is planned for a future work, but it has been explored for 1 and 2 point functions at the 1-loop level in Finn et al. (2020); Alminawi et al. (2025).